# Efficacy of Language Model Self-Play in Non-Zero-Sum Games

### v1

@@ 6.2 Dialogue Length and Diversity
We qualitatively observed that dialogues in the semi-competitive setting became less diverse over the course of self-play finetuning. We quantified this in two ways: (1) by average dialogue length, in Figure 3, and (2) by the number of unique words produced during each iteration, in Figure 4.
We first computed the average dialogue length over 500 hundred games of self-play during each iteration of model training in the semi-competitive and cooperative settings. We found that self-play caused dialogues to become substantially longer in the cooperative setting but shorter in the semi-cooperative one. We hypothesize that this discrepancy may occur because agents in the cooperative setting are incentivized to share all information; qualitatively, we often observed models sharing exact details of their private value functions. Additionally, we found that models which argue with each other for too long are more likely to “go off the rails” and fail to reach an agreement at all; because these games receive scores of zero, this behavior may be filtered out over the course of self-play under the semi-competitive objective.
We also computed the “vocabulary size” of each iteration by counting the number of unique tokens produced during 500 games of self-play. We observed a similar trend of decreasing vocabulary size over the course of self-play in the semi-competitive setting, supporting the hypothesis that the semi-competitive objective leads to convergence in model behavior. However, as shown in Section 5.2, these models performed well in both self-play and in human generalization experiments, suggesting that they may not need very diverse communication strategies to achieve high scores.
@@ 7 Related Work
Much prior work on goal-oriented dialogue has focused on collaborative settings. In tasks such as Cards (Djalali et al., 2011; Potts, 2012), CerealBar (Suhr et al., 2019), OneCommon (Udagawa and Aizawa, 2019), and DialOp (Lin et al., 2024), two agents must collaborate via natural language dialogue to achieve a shared goal within an environment. In many of these tasks, models are automatically evaluated via self-play, which serves as a proxy for human evaluation (Fried et al., 2021; Lin et al., 2024).
Relatively less work has focused on the case where agents have conflicting goals. A handful of grounded dialogue tasks are focused on bartering or negotiation, including Deal or No Deal (DoND; Lewis et al., 2017), CaSiNo (Chawla et al., 2021), and the fruit trading game from Gemp et al. (2024). These games are all structurally similar and differ primarily in the number and types of objects they use, as well as the public availability of human data. In the Craigslist Bargaining task (He et al., 2018), agents negotiate on the price of a object for sale.
Fried et al. (2023) provides additional discussion of collaborative and competitive grounded dialogue tasks and modeling approaches.
@@
Lewis et al. (2017) trained GRU-based language models on the Deal or No Deal task using REINFORCE (Williams, 1992). In contrast to our work, Lewis et al. (2017) did not learn a model tabula rasa but instead interleaved reinforcement learning from self-play with supervised learning on task-specific data to avoid divergence from human-interpretable language. Divergence issues abound in other settings where models are trained via self-play, including in emergent communication (Kottur et al., 2017; Lowe et al., 2019; Tomlin and Pavlick, 2019).
A wave of recent work has focused on methods for autonomously improving large language models at training (Ouyang et al., 2022; Bai et al., 2022; Abdulhai et al., 2023) or inference (Shinn et al., 2023; Yao et al., 2023; Wu et al., 2024) time. Chen et al. (2024) propose a method called self-play fine-tuning, but their usage of the term self-play differs from the traditional meaning, i.e., it does not involve agents interacting within an environment; instead, Chen et al. (2024) proposes a preference learning method akin to DPO (Rafailov et al., 2023) or PPO for reinforcement learning from human feedback (Ouyang et al., 2022). More closely related to our work is Pan et al. (2024), which iteratively trains models for device-control tasks using filtered behavior cloning; however, in contrast to our work, Pan et al. (2024) studies a single agent interacting with an environment, rather than multiple agents interacting with one another.
Another closely related work is Fu et al. (2023), which uses self-play to refine language models for a bargaining task. In contrast to our work, the bargaining task is a relatively simple, zero-sum game which elicits less rich dialogues than DoND. Further, while our work focuses on finetuning models, Fu et al. (2023) present models with in-context demonstrations of previous games and natural language feedback from a critic model (similar to Reflexion (Shinn et al., 2023)), leading to less major performance improvements.
@@
Training agents against copies of themselves is a longstanding technique in reinforcement learning (Littman, 1994), popularized in the past decade by models like AlphaGo (Silver et al., 2016) and AlphaZero (Silver et al., 2017). Experiments on games such as Overcooked (Carroll et al., 2019) and Hanabi (Bard et al., 2020) have shown that policies learned via self-play often fail to generalize to collaborative or imperfect information games. Methods such as fictitious self-play and population play have been proposed to address these issues (Heinrich et al., 2015; Strouse et al., 2021), but have primarily been applied in games without language components. One notable exception to this is CICERO, which trained models for the game of Diplomacy via self-play using a KL-regularization objective which prevented language from drifting too far from human-written training data (FAIR, 2022).
@@ 8 Conclusion & Discussion
Our experiments showed that language model self-play can lead to significant performance improvements in both semi-competitive and cooperative games. This finding contradicts existing wisdom that self-play is ineffective in collaborative domains (Strouse et al., 2021), or that models need to be trained on task-specific human data to avoid divergence from human-interpretable language (Lewis et al., 2017; FAIR, 2022). One hypothesis is that because we observed significant model improvements after just ten rounds of self-play, the model may not have had time to overfit to cooperation with itself. Another hypothesis is that better language models might simply be more robust to the negative effects of self-play. Given a model with good generalization abilities, finetuning on self-play games might be able to elicit model capabilities which are not directly present in the self-play data. Furthermore, self-play with pretrained language models might actually function more similarly to population play, since large language models are trained on text from a population of users and may simulate different personas in different contexts (Pataranutaporn et al., 2021; Park et al., 2023).
Although game scores increased significantly after self-play, this increase can be almost entirely attributed to an increase in the percentage of completed games, rather than better high-level strategies or negotiation tactics. Future work may be able to obtain even larger improvements by combining self-play with approaches other than filtered BC, such as natural language reflections, e.g., akin to Shinn et al. (2023). Another possible approach is described by Srivastava et al. (2024), in which a language model is used to describe distributional differences between good and bad trajectories.
Finally, the effectiveness of methods like self-play is completely dependent on a reward signal, which in this work was obtained from the game environment. To apply similar methods not just in game-playing domains but in real-world scenarios, we anticipate that models will need to rely on feedback from general-purpose, learned reward models (e.g., as in Du et al., 2023; Pan et al., 2024). We leave further investigation of the challenges associated with bringing self-play into real-world application domains to future work.
@@ Acknowledgments
We are grateful to Daniel Fried, Justin Chiu, and Jessy Lin for early discussions which contributed to the idea for this work. We also thank Lucy Li, Jiayi Pan, and Rodolfo Corona for their feedback. NT is supported by the DARPA SemaFor program.
@@ Appendix A Environment Implementation
We implemented an environment for the task under which language models can play the game with each other or against another human. For initializing a new instance of the game, we sample from a list of 4,186 valid game contexts (shared item counts and private value functions for each player), provided by Lewis et al. (2017). The environment then takes turns prompting each player to either send a message (prepended by [message]) or submit a proposal (prepended by [propose]). After detecting a submitted proposal, the environment forces the other player to submit a proposal of their own. This is enforced by error correction, described in the section below. Once a game is completed, the environment uses the game context and the submitted proposals to determine the final score of each player, conditioned on the objective under which the players were instructed to play.
@@
Outputting text without a prefix of either "[message]" or "[propose]" Your output should either begin with [message] or a [propose].
@@
Sending messages with multiple mentions of "[message]" or "[propose]" (i.e. outputting multiple messages in a row) Do not include any mentions of [message] or [propose] after the initial prefix. Please just send a single message, beginning with [message].

### v2

@@ 6.2 Dialogue Length and Diversity
One natural question is whether self-play finetuning has any adverse effects on language quality. We qualitatively observed that dialogues in the semi-competitive setting became less diverse over the course of self-play finetuning. We quantified this in two ways: (1) by average dialogue length and (2) by the number of unique words produced during each iteration, as reported in Figure 3.
We first computed the average dialogue length over 500 hundred games of self-play during each iteration of model training in the semi-competitive and cooperative settings. We found that self-play caused dialogues to become substantially longer in the cooperative setting but shorter in the semi-competitive one. We hypothesize that this discrepancy may occur because agents in the cooperative setting are incentivized to share all information; qualitatively, we often observed models sharing exact details of their private value functions. Additionally, we found that models which argue with each other for too long are more likely to “go off the rails” and fail to reach an agreement at all; because these games receive scores of zero, this behavior may be filtered out over the course of self-play under the semi-competitive objective.
We also computed the vocabulary size of each iteration by counting the number of unique word types produced during 500 games of self-play. We observed a similar trend of decreasing vocabulary size over the course of self-play in the semi-competitive setting, supporting the hypothesis that the semi-competitive objective leads to convergence in model behavior. However, as shown in Section 5.2, these models performed well in both self-play and in human generalization experiments, suggesting that they may not need very diverse communication strategies to achieve high scores. We include additional results on $n$-gram entropy in Appendix E, where we observe mostly similar trends.
6.3 Hallucinations and Grounding
Figure 4: The rate of hallucinations or otherwise inconsistent messages and proposals declines over the course of self-play finetuning. We report this value as a per-message rate, rather than per-game.
We observed that baseline models often failed to reach agreements because they hallucinated items, lied about their point values, or made proposals which contradicted their previous utterances; in contrast, self-play finetuned models rarely seemed to hallucinate. To quantify this claim, we used a stronger language model to annotate the rate at which messages or proposals in self-play dialogues exhibited inconsistencies. Specifically, we prompted GPT-4 to indicate whether each message (1) lied about the player’s point value for an item, (2) made an impossible proposal based on the item counts in the game context, or (3) made a proposal explicitly contradicting what was agreed upon in the discussion. [^2] In order to validate this method, the authors manually annotated 100 randomly selected messages across iterations and compared their predictions with those of GPT-4, obtaining 92% agreement. We found that the rate of inconsistent messages decreased from 27% to 6% in semi-competitive games and from 24% to 4% in cooperative games. In contrast to prior work, which found that self-play increased the rate of lying and deceptive behavior (Lewis et al., 2017), these results suggest that producing mostly honest and accurate messages is a reasonable strategy, even in not-fully-cooperative scenarios. We provide further details, including the prompt used to generate these results, in Appendix E.
@@ 7 Related Work
Much prior work on goal-oriented dialogue has focused on collaborative settings. In tasks such as Cards (Djalali et al., 2011; Potts, 2012), CerealBar (Suhr et al., 2019), OneCommon (Udagawa & Aizawa, 2019), and DialOp (Lin et al., 2024), two or more agents must collaborate via natural language dialogue to achieve a shared goal within an environment. In many of these tasks, models are often evaluated via self-play, which serves as a proxy for human evaluation.
Another line of work has focused on the case where agents have conflicting goals. A handful of grounded dialogue tasks are focused on bartering or negotiation, including Deal or No Deal (DoND; Lewis et al., 2017), which is based on the multi issue bargaining task from DeVault et al. (2015), as well as CaSiNo (Chawla et al., 2021) and the fruit trading game from Gemp et al. (2024). These games are all structurally similar and differ primarily in the number and types of objects they use, as well as the public availability of human data. In the Craigslist Bargaining task (He et al., 2018), agents negotiate on the price of a object for sale. Recently, several new game environments have been proposed for benchmarking language model agents (Chalamalasetti et al., 2023; Qiao et al., 2023; Li et al., 2023; Wu et al., 2024a; Gong et al., 2024). Fried et al. (2023) provides additional discussion of grounded dialogue tasks and modeling approaches.
@@
Lewis et al. (2017) trained GRU-based language models on the Deal or No Deal task using REINFORCE (Williams, 1992). In contrast to our work, Lewis et al. (2017) did not learn a model tabula rasa but instead interleaved reinforcement learning from self-play with supervised learning on task-specific data to avoid divergence from human-interpretable language. Divergence issues abound in other settings where models are trained via self-play, such as in emergent communication (Kottur et al., 2017; Lowe et al., 2019; Tomlin & Pavlick, 2019), including some work on negotiation settings (Cao et al., 2018; Noukhovitch et al., 2021).
A wave of recent work has focused on methods for autonomously improving large language models at training (Ouyang et al., 2022; Bai et al., 2022; Abdulhai et al., 2023) or inference (Shinn et al., 2023; Yao et al., 2023; Wu et al., 2024b) time. Methods like StaR (Zelikman et al., 2022) and Rest-EM (Singh et al., 2024) iteratively train models on their own filtered outputs to improve reasoning capabilities. Closely related to our work is Pan et al. (2024), which iteratively trains models for device-control tasks using filtered behavior cloning; however, in contrast to our work, Pan et al. (2024) studies a single agent interacting with an environment, rather than multiple agents interacting with one another. Another closely related paper is Fu et al. (2023), which uses self-play to refine language models for a distributive bargaining task; in contrast to our work, Fu et al. (2023) use in-context learning rather than iterative finetuning, leading to less major performance improvements. Finally, the recently proposed SOTOPIA-$\pi$ (Wang et al., 2024) trains models via a similar filtered BC method on a benchmark of social tasks (Zhou et al., 2024).
@@
Training agents against copies of themselves is a longstanding technique in reinforcement learning (Littman, 1994), popularized in the past decade by models like AlphaGo (Silver et al., 2016) and AlphaZero (Silver et al., 2017). Experiments on games such as Overcooked (Carroll et al., 2019) and Hanabi (Bard et al., 2020) have shown that policies learned via self-play often fail to generalize to collaborative or imperfect information games. Methods such as fictitious self-play and population play have been proposed to address these issues (Heinrich et al., 2015; Strouse et al., 2021), but have primarily been applied in games without language components. In games with language, a KL-regularization objective is often used to prevent language from drifting too far from human-written training data (Jaques et al., 2019; FAIR, 2022).
@@ 8 Conclusion & Discussion
Our experiments showed that language model self-play can lead to significant performance improvements in both semi-competitive and cooperative games. This finding contradicts existing wisdom that self-play is ineffective in collaborative domains (Strouse et al., 2021), or that models need to be trained on task-specific human data to avoid divergence from human-interpretable language (Lewis et al., 2017; FAIR, 2022). One hypothesis is that because we observed significant model improvements after just ten rounds of self-play, the model may not have had time to overfit to cooperation with itself. Another hypothesis is that better language models might be more robust to the negative effects of self-play. Given a model with good generalization abilities, finetuning on self-play games might be able to elicit model capabilities which are not directly present in the self-play data. Furthermore, self-play with pretrained language models might actually function more similarly to population play, since large language models are trained on text from a population of users and may simulate different personas in different contexts (Pataranutaporn et al., 2021; Park et al., 2023).
Although game scores increased significantly after self-play, this increase can be almost entirely attributed to an increase in the percentage of completed games, rather than better strategic reasoning or negotiation tactics. We anticipate that future work may be able to obtain even larger improvements by combining self-play with approaches other than filtered BC, such as natural language reflections (Shinn et al., 2023). Another possible approach is described by Srivastava et al. (2024), in which a language model is used to describe distributional differences between good and bad trajectories.
Finally, the effectiveness of methods like self-play is completely dependent on a reward signal, which in this work was obtained from the game environment. To apply similar methods in real-world settings, we anticipate that models will need to rely on feedback from general-purpose, learned evaluators (e.g., as in Du et al., 2023; Pan et al., 2024). We leave further investigation of the challenges associated with bringing self-play into real-world application domains to future work.
@@ Acknowledgments
We are grateful to Daniel Fried, Justin Chiu, and Jessy Lin for early discussions which contributed to the idea for this work. We also thank Lucy Li, Jiayi Pan, and Rodolfo Corona for their feedback. Nicholas Tomlin is supported by the DARPA SemaFor program and a grant from FAR AI.
@@ Appendix A Environment Implementation
We implemented an environment for the task under which language models can play the game with each other or against another human. For initializing a new instance of the game, we sample from a list of 4,186 valid game contexts (shared item counts and private value functions for each player), provided by Lewis et al. (2017). The environment then takes turns prompting each player to either send a message (prepended by [message]) or submit a proposal (prepended by [propose]). After detecting a submitted proposal, the environment forces the other player to submit a proposal of their own. This is enforced by error correction, as described below. Once a game is completed, the environment uses the game context and the submitted proposals to determine the final score of each player, conditioned on the objective under which the players were instructed to play.
@@
Outputting text without a prefix of either “[message]” or “[propose]” Your output should either begin with [message] or a [propose].
@@
Sending messages with multiple mentions of “[message]” or “[propose]” (i.e. outputting multiple messages in a row) Do not include any mentions of [message] or [propose] after the initial prefix. Please just send a single message, beginning with [message].
@@ Appendix B Model Training and Hyperparameters
To ensure that GPT-3.5’s low initial scores were not a result of suboptimal prompting, we manually crafted 10 alternative semi-competitive prompts (and their cooperative counterparts) and ran 100 games of self-play using GPT-3.5 for each one. For the semi-competitive setting, the mean score across prompts was 0.382, with a minimum of 0.18 and a maximum of 0.52. For the cooperative setting, the mean score across prompts was 0.636, with a minimum of 0.25 and a maximum of 1.17.
@@ Appendix D Results for Strict Competition
Appendix E Additional Analysis Details
Additional Results
We include additional error and abort rates in Figure 10. To further quantify the diversity of self-play finetuned models, we also fit $n$-gram models on the dialogues produced during each iteration of self-play and report their results in Figure 11.
Details of Hallucination Analysis
We used GPT-4 to analyze the frequency of hallucinations and inconsistent messages or proposals in self-play dialogues. Our analysis prompt, provided in Figure 9, takes in an entire dialogue with one player’s private value function masked out and returns a list of binary classifications describing whether each message or proposal was consistent with the game context or not. We ran this prompt on 500 games from each iteration of self-play, for each of the semi-competitive and cooperative objectives. For each dialogue, we ran the analysis prompt twice, once from each player’s perspective. The results of this analysis are presented in Figure 4.
This is a game where players with private value functions negotiate how to divide a shared set of items and end the dialogue by making private proposals about which items they will claim for themselves. The proposals are private and final, meaning the model should submit proposals that are consistent with what the players agreed upon. Analyze the following dialogue between LLMs and determine if the messages of the dialogue from Player {player_to_analyze} either: - lies about the player’s point value for an item, - makes an impossible proposal based on the item counts in the game context, or - makes a proposal explicitly contradicting what was agreed upon in the discussion Your output should be a structured JSON output, containing three things for each message: - "message": a repeat of the line (including the source, e.g. "Player {player_to_analyze}") - "analysis": a step-by-step analysis of whether the message violates any of the conditions above - "decision": a final [YES] or [NO] answer. Dialogue: {conversation_history}
Figure 9: Prompt used for computing the rate of hallucinations and inconsistent messages or proposals in self-play dialogues. We used GPT-4 to classify whether each message or proposal was consistent with the game context so far, finding that consistency increased over the course of self-play finetuning. Values in {brackets} are filled in based on the game context.
Figure 10: Rate of self-play games consisting of errors (left) or aborts (right) over the course of self-play finetuning. When a model produces an invalid message or proposal, it receives an error message from the environment and is given an opportunity to re-generate the message. If a model makes five errors in a row, the game aborts. Errors and aborts decline over the course of training.
(a) Semi-Competitive
(b) Cooperative
Figure 11: We fit $n$-gram models on the semi-competitive (left) and cooperative (right) self-play games and computed the entropy of their distributions. As with vocabulary size, the entropy decreased in the semi-competitive setting but stayed roughly constant in the cooperative setting.